# On the presence of a fifth force at the Galactic Center

### v3

@@ 3.4 Comparison with theoretical estimates
As stated in the previous section, the inclusion of the 1PN acceleration in the equations of motion implies an additional assumption, i.e., that any correction to the GR expression is subleading with respect to Eq. (3) and hence negligible in our fit. In order to show this, we compare our results obtained using Eq. (3) with the analytical expressions derived in the literature for some specific theories.
Alves et al. (2024) derived the 1PN acceleration in massive Brans-Dicke theory, which results in exactly the same expression as Eq. (3), as long as one identifies $\alpha=a_{1}/2$, where $a_{1}=2\varphi/(2\omega_{0}+3)$, and sets the reduced mass $\eta=0$, which is clearly a good approximation for the SgrA-S2222 system.
In Tan & Lu (2024) an analytical expression for the 1PN acceleration in $f(R)$ gravity is obtained and reported in their Eq. $(17)$. We used this expression to show that our upper limits on $\alpha$ are not affected by this difference in the 1PN expansion, at least when the uncertainty on $\alpha$ is the smallest, that is, for $\lambda=3\cdot 10^{13}\,\rm m$. In Figure 3 the comparison between the posteriors is reported, showing that the upper limit on $\alpha$ is only changed by a factor $2$ when the full expression for $f(R)$ gravity is used.
Tan & Lu (2024) also showed that the use of a multiple-star fit (specifically including S2222, S29292929 and S55555555) with the $1$PN acceleration for $f(R)$ gravity derived in Eq. $(17)$, could potentially break the degeneracy between $M$ and $\alpha$ in the large $\lambda$ limit, producing a stringent upper limit on the fifth force intensity also for $\lambda>10^{15}\,\rm m$.
In addition, in Losada et al. (2025) it is shown that combining S2222 motion with S62626262 lensing observations can also potentially break the degeneracy between the parametrized PN parameters $\beta$ and $\gamma$. We leave the multi-star analysis for future work.
Figure 3: Posterior probability density of $|\alpha|$ using the $1$PN expression in Eq. (3) (black curve) versus the $1$PN expansion for $f(R)$ gravity derived in Tan & Lu (2024) (red curve), when $\lambda=3\cdot 10^{13}$ m.
@@ 4 Conclusions
In this paper, we update the current constraints on the fifth force intensity at the GC using GRAVITY data for S2222 from $2017$ to $2022$, including the pericenter passage. Those data allowed us to significantly improve previous estimates on the same effect, giving a $95\%$ confidence level curve that is one order of magnitude below the previous estimates (see Figure 2). Specifically, three different behaviors are found in the posterior distribution of $\alpha$, according to the value of the length scale of the new interaction $\lambda$ compared to S2222 orbital range.
The minimum value of $\alpha$ is found for $\lambda=3\cdot 10^{13}\,\rm m\,(\sim 200\,AU)$ where $|\alpha|<0.003$. For comparison, in correspondence of the minimum found by Hees et al. (2017), $\lambda=150\,\rm AU\sim 2.2\cdot 10^{13}\,m$, we found $|\alpha|<0.0031$.
We also showed that the $1$PN expansion used in this work coincides with the expression developed for massive Brans-Dicke theory and that additional terms proportional to $\alpha$ in the $1$PN acceleration for $f(R)$ theories are subdominant with respect to the expression used in this work and hence negligible, as the upper limit on $|\alpha|$ is unaffected.
A complete analysis including all S-stars, specifically those with apocenter passage farther away than S2222, is left for future work, with the aim of further improving the confidence level curve, possibly expanding the range of $\lambda$ and obtaining meaningful constraints also at $\lambda\gtrsim 8\cdot 10^{14}\,\rm m$.
@@ Acknowledgements.
Appendix A Details about numerical integration
The numerical integration of the equation of motion is performed using the Python library scipy.integrate.solve_ivp with a Runge-Kutta 5(4) algorithm, which means that the steps are evaluated using a 5-th order method, while the error is controlled assuming the accuracy of the 4-th order method. The convergence of the integration is ensured by looking at the conservation of energy over the entire integration period (almost two orbits in $\sim 30$ years gives $\Delta E/E\sim\mathscr{O}(10^{-10})$).
Kepler’s equation is solved instead using a Python root finder (scipy.optimize.newton) which implements the Newton-Raphson method. The latter solves the equation with precision of $\mathscr{O}(10^{-16})$.
@@ B.1 Coordinate transformation
$\begin{split}&x^{\prime}=Ax_{\rm BH}+Fy_{\rm BH}\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,v_{x^{\prime}}=Av_{x_{\rm BH}}+Fv_{y_{\rm BH}}\\ &y^{\prime}=Bx_{\rm BH}+Gy_{\rm BH}\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,v_{y^{\prime}}=Bv_{x_{\rm BH}}+Gv_{y_{\rm BH}}\\ &z_{\rm obs}=-(Cx_{\rm BH}+Hy_{\rm BH})\,\,\,\,\,\,\,\,\,\,\,\,v_{z_{\rm obs}}=-(Cv_{x_{\rm BH}}+Hv_{y_{\rm BH}})\,,\end{split}$ (9)
@@
$\begin{split}&A=\cos\Omega\cos\omega-\sin\Omega\sin\omega\cos i\\ &B=\sin\Omega\cos\omega+\cos\Omega\sin\omega\cos i\\ &F=-\cos\Omega\sin\omega-\sin\Omega\cos\omega\cos i\\ &G=-\sin\Omega\sin\omega+\cos\Omega\cos\omega\cos i\\ &C=-\sin\omega\sin i\\ &H=-\cos\omega\sin i\,,\end{split}$ (10)
while the Cartesian coordinates $\{x_{\rm BH},y_{\rm BH},z_{\rm BH}\}$ and velocities $\{v_{x_{\rm BH}},v_{y_{\rm BH}},v_{z_{\rm BH}}\}$ are those obtained from the numerical integration. For a more detailed discussion about how the coordinate system $\{x^{\prime},y^{\prime},z_{\rm obs}\}$ and the above transformation are defined we refer the reader to Figure 1 and Appendix B of Grould et al. (2017).
@@ B.2 Relativistic effects and Rømer’s delay
In order to produce a better fit, there are observational effects that must be included in the model.
The Rømer’s delay is the difference between the time of emission of the signal $t_{\rm em}$ and the actual observational dates $t_{\rm obs}$, due to the finite speed of light. To include this delay, we used the first order Taylor’s expansion of the Røemer equation, which reads:
@@
The difference between the exact solution of Røemer equation and the approximated solution in (11) is at most $\sim 4$ s over the S2222 orbit and therefore negligible. The Rømer effect affects both the astrometry and the spectroscopy, with an impact of $\approx 450\,\mu$as on positions and $\approx 50$ km/s at periastron on radial velocities. Our results recover the previous estimates for this effect reported in Grould et al. (2017); GRAVITY Collaboration (2018).
Moreover, there are two relativistic effects that must be taken into account when S2222 approaches the periastron: the relativistic Doppler shift and the gravitational redshift. Both induce a shift in the spectral lines of S2222 that affects the radial velocity measurements. The former is given by
@@
In the total space velocity $\textbf{v}|$ v | we must also add a correction due to the Solar System motion. We followed the most recent work of Reid & Brunthaler (2020) and take a proper motion of Sgr A of

### v4

@@ 3.4 Comparison with theoretical estimates
As stated in the previous section, the inclusion of the 1PN acceleration in the equations of motion implies an additional assumption; namely, that any correction to the GR expression is subleading with respect to Eq. (3) and, hence, negligible in our fit. To show this, we compared our results obtained using Eq. (3) with the analytical expressions derived in the literature for some specific theories.
Alves et al. (2024) derived the 1PN acceleration in massive Brans-Dicke theory, which results in exactly the same expression as Eq. (3), as long as we identify $\alpha=a_{1}/2$, where $a_{1}=2\varphi/(2\omega_{0}+3)$, and set the reduced mass to $\eta=0$, which is clearly a good approximation for the SgrA-S22 system.
In Tan and Lu (2024) an analytical expression for the 1PN acceleration in $f(R)$ gravity is obtained and reported in their Eq. $(17)$. We used this expression to show that our upper limits on $\alpha$ are not affected by this difference in the 1PN expansion, at least when the uncertainty on $\alpha$ is the smallest (i.e., for $\lambda=3\cdot 10^{13}\,\rm m)$. In Fig. 3, the comparison between the posteriors is reported, showing that the upper limit on $\alpha$ is only changed by a factor of $2$ when the full expression for $f(R)$ gravity is used.
Tan and Lu (2024) also showed that the use of a multiple-star fit (specifically including S22, S29,29, and S5555) with the 11PN acceleration for $f(R)$ gravity derived in Eq. $(17)$ could potentially break the degeneracy between $M$ and $\alpha$ in the large $\lambda$ limit. This could produce a stringent upper limit on the fifth force intensity for $\lambda>10^{15}\,\rm m$ as well.
In addition, Losada et al. (2025) demonstrated that combining S22 motion with S6262 lensing observations can also potentially break the degeneracy between the parametrized PN parameters $\beta$ and $\gamma$. However, we leave a multi-star analysis for a future work.
Figure 3: Posterior probability density of $|\alpha|$ using the 11PN expression in Eq. (3) (black curve) versus the 11PN expansion for $f(R)$ gravity derived in Tan and Lu (2024) (red curve), when $\lambda=3\cdot 10^{13}$ m.
@@ 4 Conclusions
In this paper, we update the current constraints on the fifth force intensity at the GC using GRAVITY data for S22 from $2017$ to $2022$, including the pericenter passage. These data have allowed us to significantly improve previous estimates on the same effect, giving a $95\%$ confidence level curve that is one order of magnitude below the previous estimates (see Figure 2). Specifically, three different behaviors are found in the posterior distribution of $\alpha$, according to the value of the length scale of the new interaction $\lambda$ compared to S22 orbital range.
The minimum value of $\alpha$ is found for $\lambda=3\cdot 10^{13}\,\rm m\,(\sim 200\,AU)$ where $|\alpha|<0.003$. For comparison, in correspondence with the minimum found by Hees and others (2017), $\lambda=150\,\rm AU\sim 2.2\cdot 10^{13}\,m$, we found $|\alpha|<0.0031$. We also showed that the 11PN expansion used in this work coincides with the expression developed for massive Brans-Dicke theory and that additional terms proportional to $\alpha$ in the 11PN acceleration for $f(R)$ theories are subdominant with respect to the expression used in this work and thereby negligible, as the upper limit on $|\alpha|$ is unaffected.
A complete analysis including all S-stars, specifically those with apocenter passage farther away than S22, is left for future work, with the aim of further improving the confidence level curve. This would allow us to possibly expand the range of $\lambda$ and obtain additional and meaningful constraints at $\lambda\gtrsim 8\cdot 10^{14}\,\rm m$.
@@ Acknowledgements.
Appendix A Details of the numerical integration
The numerical integration of the equation of motion is performed using the Python library scipy.integrate.solve_ivp with a Runge-Kutta 5(4) algorithm, which means that the steps are evaluated using a fifth-order method, while the error is controlled assuming the accuracy of the fourth-order method. The convergence of the integration is ensured by looking at the conservation of energy over the entire integration period (almost two orbits in $\sim 30$ years gives $\Delta E/E\sim\mathscr{O}(10^{-10})$).
Kepler’s equation can be solved instead using a Python root finder (scipy.optimize.newton), which implements the Newton-Raphson method. The latter solves the equation with a precision of $\mathscr{O}(10^{-16})$.
@@ B.1 Coordinate transformation
[…]
@@
$\begin{split}&A=\cos\Omega\cos\omega-\sin\Omega\sin\omega\cos i,\\ &B=\sin\Omega\cos\omega+\cos\Omega\sin\omega\cos i,\\ &F=-\cos\Omega\sin\omega-\sin\Omega\cos\omega\cos i,\\ &G=-\sin\Omega\sin\omega+\cos\Omega\cos\omega\cos i,\\ &C=-\sin\omega\sin i,\\ &H=-\cos\omega\sin i\,.\end{split}$ (10)
The Cartesian coordinates $\{x_{\rm BH},y_{\rm BH},z_{\rm BH}\}$ and velocities $\{v_{x_{\rm BH}},v_{y_{\rm BH}},v_{z_{\rm BH}}\}$ are those obtained from the numerical integration. For a more detailed discussion on how the coordinate system $\{x^{\prime},y^{\prime},z_{\rm obs}\}$ and the above transformation are defined, we refer to Figure 1 and Appendix B of Grould et al. (2017).
@@ B.2 Relativistic effects and Rømer’s delay
To produce a better fit, there are observational effects that must be included in the model.
Rømer’s delay is the difference between the time of emission of the signal $t_{\rm em}$ and the actual observational dates $t_{\rm obs}$, due to the finite speed of light. To include this delay, we used the first order Taylor’s expansion of the Røemer equation, which is expressed as:
@@
[…]
Moreover, there are two relativistic effects that must be taken into account when S22 approaches the periastron: the relativistic Doppler shift and the gravitational redshift. Both induce a shift in the spectral lines of S22 that affects the radial velocity measurements. The former is given by
@@
In the total space velocity, $v=|\textbf{v}|$, we must also add a correction due to Solar System motion. We followed the most recent work of Reid and Brunthaler (2020) and take a proper motion of Sgr A of